OpenStudy (anonymous):

The table shows the annual consumption of cheese per person in the United States for selected years in the 20th century. Let x = number of years since 1900, and y = pounds per person. What cubic model best fits this data? Year Pounds Consumed 1902 1.959 1924 6.373 1946 7.29 1982 15.817

OpenStudy (anonymous):

A. y = 0.00989x3 - 0.0000872x2 + 1.192x + 0.403 B. y = -0.0000872x3 + 0.00989x2 - 0.403x - 1.192 C. y = -0.0000872x3 - 0.00989x2 - 0.403x + 1.192 D. y = 0.0000872x3 - 0.00989x2 + 0.403x + 1.192